OpenStudy (anonymous):

factor 2x^2 − 8x − 42. @amistre64

OpenStudy (anonymous):

algebra .... simplify it

OpenStudy (anonymous):

How... ?

OpenStudy (luigi0210):

Start by factoring that 2 out.

OpenStudy (anonymous):

s divide by 2

OpenStudy (anonymous):

x^2-4x-21

OpenStudy (anonymous):

x^2+3x-7x-21 =x(x+3)-7(x+3) =(x-7)(x+3) therefore x=7,-3

OpenStudy (anonymous):

be careful with the sign .... .....+3-7=-4 did u understand ??

OpenStudy (mathstudent55):

2x^2 − 8x − 42 Factor out 2: 2(x^2 - 4x - 21) What two numbers multiply to -21 and add to -4? -7 and 3, so 2(x - 7)(x + 3) The problem only asks for factoring, so keep the 2 as a factor.